### v2

@@ Abstract
Accurately predicting blood glucose (BG) levels of ICU patients is critical, as both hypoglycemia (BG < 70 mg/dL) and hyperglycemia (BG > 180 mg/dL) are associated with increased morbidity and mortality. This study presents a proof-of-concept machine learning framework, the Multi-source Irregular Time-Series Transformer (MITST), designed to predict blood glucose (BG) levels in ICU patients. Unlike existing approaches that rely on manual feature engineering or are limited to a small number of Electronic Health Record (EHR) data sources, MITST demonstrates the feasibility of integrating diverse clinical data (e.g., lab results, medications, vital signs) and handling irregular time-series data without predefined aggregation. MITST employs a hierarchical architecture of Transformers, comprising feature-level, timestamp-level, and source-level components, to capture fine-grained temporal dynamics and enable learning-based data integration. This eliminates the need for traditional aggregation and manual feature engineering. In a large-scale evaluation using the eICU database (200,859 ICU stays across 208 hospitals), MITST achieves an average improvement of 1.7% ($p<0.001$) in AUROC and 1.8% ($p<0.001$) in AUPRC over a state-of-the-art baseline. For hypoglycemia, MITST achieves an AUROC of 0.915 and an AUPRC of 0.247, both significantly outperforming the baseline. The flexible architecture of MITST allows seamless integration of new data sources without retraining the entire model, enhancing its adaptability for clinical decision support. While this study focuses on predicting BG levels, MITST can easily be extended to other critical event prediction tasks in ICU settings, offering a robust solution for analyzing complex, multi-source, irregular time-series data.
@@ 1 Introduction
Managing BG levels in ICU patients is particularly challenging, as both hypoglycemia and hyperglycemia are associated with increased length of stay, morbidity, and mortality [8, 9, 10, 11]. Hypoglycemia is a condition characterized by abnormally low BG levels, below 70 mg/dL, while hyperglycemia refers to abnormally high BG levels, above 180 mg/dL [12]. Especially people with diabetes are prone to hypoglycemia and hyperglycemia due to their impaired ability to regulate BG levels effectively. The prevalence of type 2 diabetes (T2DM) has reached alarming proportions globally, affecting approximately 462 million individuals worldwide, and posing a significant burden on healthcare systems [13]. Individuals with T2DM are at heightened risk of dysglycemia in the ICU. These acute glucose fluctuations are associated with increased morbidity and mortality rates, particularly in critical care settings [14]. Effective prediction of dysglycemic events is therefore essential to mitigate adverse outcomes, yet the dynamic nature of glycemic control in ICU patients makes this prediction highly challenging.
@@
Overcoming these challenges could enhance the accuracy, interoperability, and clinical relevance of EHR data analysis, ultimately improving healthcare outcomes. We propose MITST (an abbreviation for Multi-source Irregular Time-Series Transformer), a novel method designed to address the irregular nature of EHR data, while preserving essential temporal details often lost with conventional approaches. MITST integrates multiple sources of medical information, creating shared representations of medical concepts and entities. Our framework provides a more robust and insightful analysis of irregular time-series forecasting with a focus on timely BG level prediction tasks, improving patient outcomes in critical care settings. Our contributions are summarized as follows.
@@
  • In our experimental results, MITST outperformed the state-of-the-art baseline model [19] for predicting BG levels, demonstrating a 1.7% improvement in the average AUROC and a 1.8% improvement in the average AUPRC across all classes. The gains were even more pronounced for clinically significant categories, particularly hypoglycemia and hyperglycemia, where accurate predictions are critical for patient outcomes. These findings underscore the superior predictive accuracy and robustness of MITST in managing complex, irregular time-series data from ICU settings.
@@ 2 Related work
BG prediction methods can be broadly classified into two categories: physiology-based and data-driven approaches [21, 22, 23]. Physiology-based methods model BG dynamics using compartments such as meal absorption, insulin activity, exercise, and individual glucose metabolism [21]. However, our focus is on data-driven approaches. For a detailed review of physiology-based methods, we refer interested readers to [23]. Data-driven methods for BG prediction fall under the broader category of clinical time-series forecasting. These datasets are often irregularly sampled, making traditional time-series models unsuitable without pre-processing steps like data imputation or window-based aggregation. Imputation-based techniques [24] aim to convert irregular time-series into regular ones for use with standard forecasting models. For example, Lipton et al. [25] applied forward- and back-filling methods on 13 clinical time-series features and used long short-term memory (LSTM) networks to predict diagnoses. Similarly, Harutyunyan et al. [26] combined imputation with indicator variables and evaluated an LSTM on four clinical tasks. Alternatively, window-based approaches aggregate task-specific features from time-series segments. For example, Harutyunyan et al. [26] used windowed features with logistic regression for ICU mortality prediction.
@@
Some studies have attempted to directly model ITS data by learning a continuous-time model from discrete and irregular observations [27, 28]. For example, [28] models ITS using multiple Gaussian processes (GPs) in the lower level of a hierarchical system and captures the transitions between GPs by using a linear dynamical system. With advancements in deep learning, ITS-focused methods have emerged. Baytas et al. [29] proposed Time-Aware LSTM (T-LSTM), which adjusts hidden states based on time gaps. Bai et al. [30] used an attention mechanism to handle irregular time-intervals between clinical events for diagnosis prediction. Tipirneni et al. [31] introduced a Transformer-based framework for ITS, learning triplet embeddings (feature name, value, timestamp) to predict mortality.
@@
Unlike most existing methods, MITST inherently handles ITS data without requiring time-series-level imputation. Instead, imputation is applied only at the feature level to fill missing values within timestamp-specific vectors, preserving the raw irregularity of the time-series data. Two notable exceptions are T-LSTM [29] and the Transformer-based model by Tipirneni et al. [31], which also adapt to the irregularity of time intervals directly within their architectures. T-LSTM models patient data as a sequence of discrete (i.e., non-consecutive) visits, requiring larger datasets across multiple patient encounters, whereas MITST focuses on data from a single ICU stay, making it applicable to scenarios with fewer data points. Also, while Tipirneni et al. evaluated their Transformer-based model on mortality prediction—a simpler task compared to next BG level forecasting—MITST is tested on a more complex task and eliminates the need for manual feature selection, which was required in their approach. Moreover, while most prior works use a limited number of EHR sources (e.g., diagnosis or procedure codes), MITST integrates a broader range of clinical data (e.g., vital signs, lab results, medications). This allows the model itself to determine the the most relevant combination of data sources for a specific clinical task, eliminating the need for predefined feature engineering that is common in traditional ITS methods.
@@
It is worth noting that Transformers, though dominant in natural language processing, are less mature in time-series forecasting [32]. Traditional machine learning models, such as random forests, XGBoost, or ARIMA combined with engineered features, remain competitive in this domain. This also holds for large language models (LLMs) in medical applications [33]. Thus, we compare MITST to the state-of-the-art random forest model developed by Zale et al. [19], which employs carefully designed features for next BG prediction.
@@ 3 Problem statement
The inputs consist of ITS data from $M$ distinct sources (e.g., medication records, lab test results, vital sign measurements). Each source provides data recorded at varying, often irregular, intervals. The training dataset consists of several labeled examples of the form $(\mathbf{x},\mathbf{y})$, where $\mathbf{y}\in\{0,1\}^{C}$ is a one-hot encoded label vector representing the class of the next BG measurement. The total number of classes is denoted by $C$ (i.e., categories for the levels of the next BG measurement - hypoglycemia, hyperglycemia, or euglycemia). Notice that the next BG measurements have been conducted at different time intervals between 5 minutes up to 10 hours for various labeled examples in our dataset.
@@
During inference time, given a new data point, the output of the model is a vector of logits $\hat{\mathbf{y}}\in\mathbb{R}^{C}$, where each element of the vector represents the raw, un-normalized score for each class before applying a softmax function to obtain the final predicted class label. The prediction task aims to estimate the class of the next BG measurement based on the data available up to the current measurement.
@@ 4 Data material
To maintain data continuity, consecutive ICU stays for the same patient are consolidated into a single stay when applicable. We include only ICU stays with at least six BG measurements, ensuring that for each input data point, there are at least five prior BG measurements available. Furthermore, each BG measurement must be followed by a subsequent reading within a window of 5 minutes to 10 hours. This interval is chosen in accordance with thresholds established in [19], ensuring consistency with prior work while allowing our model to operate within clinically relevant prediction windows. The final dataset comprised 97,383 ICU stays, which were split across patients into three non-overlapping sets: 70% for training, 10% for validation, and 20% for testing. Fig. 1 provides a flowchart detailing the inclusion criteria, the number of unique stays, patients, and examples within each set. It also presents the distribution of examples by BG categories and includes the statistics of diabetic patients, along with the number of examples for type 1 and type 2 diabetes (T1DM and T2DM, respectively). The classification of T1DM and T2DM was determined using ICD-9 and ICD-10 codes extracted from the diagnosis table. We did our best throughout the paper to adhere to the MI-CLAIM checklist [35] to ensure the transparency and reproducibility of our study. For further details regarding the dataset and its demographics, refer to the original publication [34]. The details on preprocessing and data imputation can be found in Section 6.1.2.
@@ 5 Our method: MITST
The time-series for each source $m$ and for every data point $\mathbf{x}$, i.e., $\mathbf{x}(m,:)$, is first passed through a source-specific module (shown as “Module 1” to “Module $M$” in Fig. 2(a), respectively), where representations (or embeddings) are extracted based on the characteristics of that particular data source. These source-specific representations are then transformed to fit into a common framework, allowing the model to analyze and learn from all the data sources simultaneously. To achieve this, each source-specific representation is transformed or “projected” (via “Proj 1” to “Proj $M$” in Fig. 2(a), respectively) into a shared space of a fixed dimension. This step ensures that data from different sources, which may have different formats and features, can be compared and combined effectively within the model. Next, these unified representations are passed through a Transformer-like encoder architecture (shown as “Standard Transformer” in Fig. 2(a), utilizing multi-head attention mechanisms. This allows the model to focus on different aspects of the input data concurrently using multiple attention heads. Each head captures various patterns or relationships across the sources, leading to a more comprehensive understanding of the input information. The outputs from these attention heads are concatenated to form a unified view of each source.
Following this, a self-attentive fusion network [36, 37] is used to combine the information from all the sources. This network employs a technique called “self-attention”, which allows the model to weigh the importance of different sources relative to each other. In simpler terms, it helps the model decide which sources are most relevant when making a prediction, based on how they interact with each other. This fusion step ensures that the final representation takes into account the most critical information from all sources. This final representation is then passed to the prediction layer.
@@
The source-specific module is further elaborated in Fig. 2(b). In our overall end-to-end framework, information is hierarchically aggregated using three types of Transformers in a supervised manner across different levels.
@@
  • Feature level: An FT-Transformer [38] (shown in Fig. 2(b) combines categorical and numerical features at each time point within each source. For instance, a single time point from the medication source includes categorical features (e.g., drug name, frequency, route of administration) and numerical features (e.g., dosage). The FT-Transformer generates a unified vector representation for them at each time point, which is subsequently used in the next layers.
@@
  • Timestamp level: A standard transformer (shown as “ITS-Transformer” in Fig. 2(b) aggregates information across irregular timestamps during patient admission. This layer learns interactions between features of a source over time, providing a source-specific representation for each patient admission. Recall that “ITS” stands for “irregular time series”. We refer to this module as the ITS-Transformer to distinguish it from the other modules in two key ways. First, its functionality differs in that it aggregates across time points. Second, it incorporates time encoding (similar to positional encoding in a standard Transformer), unlike the other two layers of aggregation.
@@
  • Source level: The aggregated representations from all sources are combined using another Transformer (shown as “Standard Transformer” in Fig. 2(b), with a final self-attention network to make predictions.
@@ 5.1.2 Timestamp Aggregation (ITS-Transformer)
After obtaining the feature-level representations $\mathbf{f}(m,t_{i})$ for each time point $i$, they are aggregated across time to capture the temporal dynamics within the source $m$. We employ a standard Transformer that processes the sequence of feature embeddings $\{\mathbf{f}(m,t_{i})\}_{i=1}^{T_{m}}$. The Transformer’s temporal encoding mechanism leverages the timestamps $t_{i}$ associated with each feature embedding $\mathbf{f}(m,t_{i})$, thereby preserving the temporal order and intervals of the timestamps. This is crucial for incorporating the irregularity of time-series data. The timestamps $t_{i}$ are encoded and added to the corresponding feature embeddings $\mathbf{f}(m,t_{i})$, ensuring that the temporal aspect is fully integrated into the learning process.
@@ 5.4 Self-attentive fusion network
Here, $\mathbf{w}_{f}\in\mathbb{R}^{d_{f}\times d_{j}}$ and $\mathbf{b}_{f}\in\mathbb{R}^{d_{f}}$ are learnable parameters of the affine transformation, and $\mathbf{u}_{f}\in\mathbb{R}^{d_{f}}$ represents the context query vector that is learned during training. The attention scores $\alpha_{m}$ reflect the importance of each source. These scores are used to produce the final representation $\mathbf{v}$ by taking a weighted sum of the source-specific representations.

### v3

@@ Abstract
Accurately predicting blood glucose (BG) levels of ICU patients is critical, as both hypoglycemia (BG < 70 mg/dL) and hyperglycemia (BG > 180 mg/dL) are associated with increased morbidity and mortality. This study presents a proof-of-concept machine learning framework, the Multi-source Irregular Time-Series Transformer (MITST), designed to predict BG levels in ICU patients. In contrast to existing methods that rely heavily on manual feature engineering or utilize limited Electronic Health Record (EHR) data sources, MITST integrates diverse clinical data—including laboratory results, medications, and vital signs—without predefined aggregation. The model leverages a hierarchical Transformer architecture, designed to capture interactions among features within individual timestamps, temporal dependencies across different timestamps, and semantic relationships across multiple data sources. Evaluated using the extensive eICU database (200,859 ICU stays across 208 hospitals), MITST achieves a statistically significant ($p<0.001$) average improvement of 1.7 percentage points (pp) in AUROC and 1.8 pp in AUPRC over a state-of-the-art random forest baseline. Crucially, for hypoglycemia—a rare but life-threatening condition—MITST increases sensitivity by 7.2 pp, potentially enabling hundreds of earlier interventions across ICU populations. The flexible architecture of MITST allows seamless integration of new data sources without retraining the entire model, enhancing its adaptability for clinical decision support. While this study focuses on predicting BG levels, we also demonstrate MITST’s ability to generalize to a distinct clinical task (in-hospital mortality prediction), highlighting its potential for broader applicability in ICU settings. MITST thus offers a robust and extensible solution for analyzing complex, multi-source, irregular time-series data.
@@ 1 Introduction
Hypoglycemia, characterized by BG levels below 70 mg/dL, and hyperglycemia, defined as BG levels exceeding 180 mg/dL, are both critically important in clinical management due to their association with increased morbidity and mortality [8, 9, 10, 11, 12]. Patients with diabetes are particularly vulnerable to these glucose fluctuations because of impaired glycemic regulation. The global prevalence of type 2 diabetes mellitus (T2DM) is alarmingly high, affecting approximately 462 million individuals worldwide and imposing substantial burdens on healthcare systems [13].
@@
Overcoming these challenges could enhance the accuracy, interoperability, and clinical relevance of EHR data analysis, ultimately improving healthcare outcomes. To this end, we propose MITST (an abbreviation for Multi-source Irregular Time-Series Transformer), a novel method designed to address the irregular nature of EHR data, while preserving essential temporal details often lost with conventional approaches. MITST integrates multiple sources of medical information, creating shared representations of medical concepts and entities. Our framework provides a more robust and insightful analysis of irregular time-series forecasting with a focus on timely BG level prediction tasks, improving patient outcomes in critical care settings. Our contributions are summarized as follows.
@@
  • In our experimental results, MITST outperformed the state-of-the-art baseline model [18] for predicting BG levels, demonstrating a 1.7 percentage points (pp) improvement in the average AUROC and a 1.8 pp improvement in the average AUPRC across all classes. The gains were even more pronounced for hypoglycemia (+5.3 pp AUROC, +3.9 pp AUPRC), the most clinically significant category, where accurate predictions are critical for patient outcomes. These findings underscore the superior predictive accuracy and robustness of MITST in managing complex, irregular time-series data from ICU settings.
@@ 2 Related work
BG prediction methods can be broadly classified into two categories: physiology-based and data-driven approaches [20, 21, 22]. Physiology-based methods model BG dynamics using compartments such as meal absorption, insulin activity, exercise, and individual glucose metabolism [20]. However, our focus is on data-driven approaches. For a detailed review of physiology-based methods, we refer interested readers to [22].
@@
Data-driven methods for BG prediction fall under the broader category of clinical time-series forecasting. These datasets are often irregularly sampled, making traditional time-series models unsuitable without pre-processing steps like data imputation or window-based aggregation. Imputation-based techniques [23] aim to convert irregular time-series into regular ones for use with standard forecasting models. For example, Lipton et al. [24] applied forward- and back-filling methods on 13 clinical time-series features and used long short-term memory (LSTM) networks to predict diagnoses. Similarly, Harutyunyan et al. [25] combined imputation with indicator variables and evaluated an LSTM on four clinical tasks. Alternatively, window-based approaches aggregate task-specific features from time-series segments. For example, Harutyunyan et al. [25] used windowed features with logistic regression for ICU mortality prediction.
@@
Some studies have attempted to directly model ITS data by learning a continuous-time model from discrete and irregular observations [26, 27]. For example, [27] models ITS using multiple Gaussian processes (GPs) in the lower level of a hierarchical system and captures the transitions between GPs by using a linear dynamical system. With advancements in deep learning, ITS-focused methods have emerged. Baytas et al. [28] proposed Time-Aware LSTM (T-LSTM), which adjusts hidden states based on time gaps between observations. Bai et al. [29] used an attention mechanism to handle irregular time-intervals between clinical events for diagnosis prediction. Tipirneni et al. [30] introduced STraTS, a Transformer-based framework for ITS, which learns triplet embeddings (feature name, value, timestamp) to predict mortality.
@@
Unlike most existing methods, MITST inherently handles ITS data without requiring time-series-level imputation. Instead, imputation is applied only at the feature level to fill missing values within timestamp-specific vectors, preserving the raw irregularity of the time-series data. Two notable exceptions that also adapt to the irregularity of time intervals directly within their architectures are T-LSTM [28] and STraTS [30]. T-LSTM models patient data as a sequence of discrete (i.e., non-consecutive) visits, requiring larger datasets across multiple patient encounters. In contrast, MITST focuses on data from a single ICU stay, making it applicable to scenarios with fewer data points. STraTS [30] handles broader data sources by modeling clinical events as triplets. However, while STraTS supports continuous-valued features, it lacks built-in support for sources with categorical features such as diagnosis codes and treatment types, which limits its utility for comprehensive EHR modeling. In contrast, MITST is designed to jointly embed heterogeneous sources—including both numerical and categorical data—with minimal feature engineering. Additionally, STraTS evaluated their model on mortality prediction, which is a simpler task compared to next BG level forecasting. Furthermore, most prior works use a limited number of EHR sources (e.g., diagnosis or procedure codes), whereas MITST integrates a broader range of clinical data (e.g., vital signs, lab results, medications). This allows MITST to determine the most relevant combination of data sources for a specific clinical task, eliminating the need for predefined aggregation common in traditional ITS methods.
@@
It is worth noting that Transformers, though dominant in natural language processing, are less mature in time-series forecasting [31]. Traditional machine learning models, such as random forests, XGBoost, or ARIMA combined with handcrafted features, remain competitive in this domain. This also holds for large language models (LLMs) in medical applications [32]. Thus, we compare MITST to the state-of-the-art random forest model developed by Zale et al. [18], which employs carefully designed features for next BG prediction.
@@ 3 Problem statement
The inputs consist of ITS data from $M$ distinct sources (e.g., medication records, lab test results, vital sign measurements). Each source provides data recorded at varying, often irregular, intervals. The training dataset consists of several labeled examples of the form $(\mathbf{x},\mathbf{y})$, where $\mathbf{y}\in\{0,1\}^{C}$ is a one-hot encoded label vector representing the class of the next BG measurement. The total number of classes is denoted by $C$ (i.e., categories for the levels of the next BG measurement - hypoglycemia, hyperglycemia, or euglycemia). Note that the intervals between consecutive BG measurements vary across the dataset, ranging from 5 minutes to 10 hours.
@@
At inference time, given a new input, the model outputs a logit vector $\hat{\mathbf{y}}\in\mathbb{R}^{C}$, where each element represents the raw, unnormalized score for each class. A softmax function is then applied to produce the final predicted label. The objective is to predict the class of the next BG measurement using only the data available up to the current BG measurement.
@@ 4 Data material
To maintain temporal continuity, we consolidated ICU stays for the same patient only when the admissions occurred at the same hospital and overlapped in time, such as transfers from the hospital floor to ICU. Specifically, we used the “unitVisitNumber” column from the eICU patient table to identify sequential visits and retained only those stays that shared identical values for “hospitalAdmitTime24”, “hospitalDischargeTime24”, and “uniquepid” columns. This ensured that the merged stays reflected a single continuous clinical episode. All timestamps were recalculated relative to the adjusted “hospitaladmitoffset”, which was reset to zero for each consolidated record. This preprocessing step ensured that the time axis remained consistent and prevented label leakage across merged segments. We include only ICU stays with at least six BG measurements, ensuring that for each input data point, there are at least five prior BG measurements available. Furthermore, each BG measurement must be followed by a subsequent reading within a window of 5 minutes to 10 hours. This interval is chosen in accordance with thresholds established in [18], ensuring consistency with prior work while allowing our model to operate within clinically relevant prediction windows. The final dataset comprised 97,383 ICU stays, which were split across patients into three non-overlapping sets: 70% for training, 10% for validation, and 20% for testing. Fig. 1 provides a flowchart detailing the inclusion criteria, the number of unique stays, patients, and examples within each set. It also presents the distribution of examples by BG category and includes the statistics of diabetic patients, along with the number of examples for type 1 and type 2 diabetes (T1DM and T2DM, respectively). The classification of T1DM and T2DM was determined using ICD-9 and ICD-10 codes extracted from the diagnosis table. We did our best throughout the paper to adhere to the MI-CLAIM checklist [34] to ensure the transparency and reproducibility of our study. For further details regarding the dataset and its demographics, refer to the original publication [33]. The details on preprocessing and data imputation can be found in Section 6.1.2.
@@ 5 Our method: MITST
The time-series for each source $m$ and for every data point $\mathbf{x}$, i.e., $\mathbf{x}(m,:)$, is first passed through a source-specific module (shown as “Module 1” to “Module $M$” in Fig. 2(a), respectively), where representations (or embeddings) are extracted based on the characteristics of that particular data source. These source-specific representations are then projected into a shared latent space (via “Proj 1” to “Proj $M$” in Fig. 2(a)) enabling heterogeneous sources to be jointly analyzed despite differences in format and dimensionality.
Next, these unified representations are passed through a Transformer-based encoder architecture (shown as “Standard Transformer” in Fig. 2(a)), utilizing multi-head attention mechanisms. Each attention head captures distinct inter-source patterns, allowing the model to construct a comprehensive contextual representation. Subsequently, a self-attentive fusion network [35, 36] integrates the source-specific embeddings by adaptively weighting their contributions. This mechanism enables the model to prioritize sources based on their contextual relevance to the prediction task. This final representation is then passed to the prediction layer.
@@
The source-specific module is further elaborated in Fig. 2(b). In our overall end-to-end framework, information is hierarchically aggregated using three Transformer-based modules, each operating at a distinct level:
@@
  • Feature level: An FT-Transformer [37] (shown in Fig. 2(b)) combines categorical and numerical features at each time point within each source. For instance, a single time point from the medication source includes categorical features (e.g., drug name, frequency, route of administration) and numerical features (e.g., dosage). The FT-Transformer generates a unified vector representation for them at each time point, which is subsequently used in the next layers.
@@
  • Timestamp level: A standard transformer (shown as “ITS-Transformer” in Fig. 2(b)) aggregates information across irregular timestamps during patient admission. This layer learns interactions between features of a source over time, providing a source-specific representation for each patient admission. Recall that “ITS” stands for “irregular time series”. We refer to this module as the ITS-Transformer to distinguish it from the other modules in two key ways. First, its functionality differs in that it aggregates across time points. Second, it incorporates time encoding (similar to positional encoding in a standard Transformer), unlike the other two layers of aggregation.
@@
  • Source level: The aggregated representations from all sources are combined using another Transformer (shown as “Standard Transformer” in Fig. 2(b)), followed by a final self-attention network to make predictions.
@@ 5.1.2 Timestamp Aggregation (ITS-Transformer)
After obtaining the feature-level representations $\mathbf{f}(m,t_{i})$ for each time point $i$, they are aggregated across time to capture the temporal dynamics within the source $m$. We employ a standard Transformer that processes the sequence of feature embeddings $\{\mathbf{f}(m,t_{i})\}_{i=1}^{T_{m}}$. The timestamps $t_{i}$ are encoded and added to the corresponding feature embeddings $\mathbf{f}(m,t_{i})$ using sinusoidal positional encodings based on their minute-level offset from ICU admission, ensuring that the temporal aspect is fully integrated into the learning process. To handle variable sampling frequencies, we do not resample or aggregate sources into fixed intervals. Instead, we retain the original sampling rates and pass the minute-level timestamps directly into each source-specific transformer. This design preserves temporal granularity, especially for sparse sources such as laboratory tests. Note that no explicit weighting across sources is introduced at this stage.
@@ 5.4 Self-attentive fusion network
Here, $\mathbf{w}_{f}\in\mathbb{R}^{d_{f}\times d_{j}}$ and $\mathbf{b}_{f}\in\mathbb{R}^{d_{f}}$ are learnable parameters of the affine transformation, and $\mathbf{u}_{f}\in\mathbb{R}^{d_{f}}$ represents the context query vector that is learned during training. The attention scores $\alpha_{m}$ reflect the contextual relevance of each source. These scores are used to produce the final representation $\mathbf{v}$ by taking a weighted sum of the source-specific representations.
This layer learns source-specific attention weights via context-aware interaction between each source embedding and a trainable query vector. This data-driven mechanism adaptively down-weights noisy or redundant signals from densely sampled sources (e.g., vitals) when sparse sources (e.g., labs) provide more informative cues for the prediction task, without requiring handcrafted rules or pre-defined weighting for each source. For example, during hypoglycemia prediction, lab measurements such as blood glucose or serum insulin may receive higher weights than vital signs, which fluctuate more frequently but are less predictive in this context.